OpenStudy (anonymous):

Juwan is 7 years older than Molly. In 5 years the sum of their ages will be 53. How old is Juwan now?

OpenStudy (phantomcrow):

Do you have an idea of what type of problem this may be?

OpenStudy (anonymous):

I believe it is an application problem using algebra

OpenStudy (phantomcrow):

Yes! Do you know what specific concept of algebra?

OpenStudy (anonymous):

Well here's the information I got with it; Solve application problems involving consecutive integers, geometry, percents, and uniform motion with one unknown.

OpenStudy (phantomcrow):

The way I solved it is by setting up a system of equations.

OpenStudy (anonymous):

Okay so how would I do that?

OpenStudy (phantomcrow):

\[J=M+7, (J+5)+(M+5)=53\]

OpenStudy (phantomcrow):

Because Juwan is seven years older that Monica, hence J=M+7 and the sum of their ages will be 53 in 5 years (5 plus their current age). Now, you may use the substitution method to solve for their age.

OpenStudy (anonymous):

Okay what is the substitution method?

OpenStudy (anonymous):

j+5+m+5=53 if j=m+7 then the substitution method is to put m+7 in j's place so m+7+5+m+5=53 2m+17=53 -17 -17 2m=36 m=18 right now m=18 and j=25

OpenStudy (anonymous):

Juwan is 25 years old.

OpenStudy (anonymous):

The substitution method basically means that if two things are equal, you can swap them. For example, if the equation is b=a+2 and you knew that b=3, you could change the problem to 3=a+2 in order to solve for a=1. Does that make sense?

OpenStudy (anonymous):

I get it now, thank you so much